OpenStudy (anonymous):

Which of the following is an identity? A. Cosx(2sinx+1) = 0 B. sin^2 x = 4 - 2cos^2 x C. (tanx + cotx)(sinx cosx) = 1 D. secx tanx - cosx cotx = sinx

OpenStudy (anonymous):

\[\tan(x) =\frac{ \sin(x) }{ \cos(x) } \]

OpenStudy (anonymous):

\[\cot(x) = \frac{ \cos(x) }{ \sin(x)}\]

OpenStudy (anonymous):

what are you doing ._.

OpenStudy (anonymous):

\[(\frac{ \sin(x) }{ \cos(x) } + \frac{ \cos(x) }{ \sin(x) })(\sin(x)\cos(x))=1\]

OpenStudy (anonymous):

????? is that the answer or what

OpenStudy (anonymous):

\[= \sin^2(x) + \cos^2(x)\]

OpenStudy (anonymous):

if you aren't ging to explain stop talking seriously...its to late for somone to do that.

OpenStudy (anonymous):

to solve this problem you need to know trig identities, so have them handy or memorized.

OpenStudy (anonymous):

an identity is an equation that is the same no matter what value u put in.

OpenStudy (anonymous):

what I did was I looked at each of the answers and manipulated them to see if I could make them look like one of the trig identities, so change tan(x) to sin(x)/cos(x) or change sec(x) to 1/cos(x) etc

OpenStudy (anonymous):

so I did this and found that i could manipulate answer c to look like this \[\sin^2(x) + \cos^2(x) = 1\]

OpenStudy (anonymous):

which is a trig identity

OpenStudy (anonymous):

so to solve this problem you need to play around with each answer until you see a trig identity